OpenStudy (anonymous):

how many different triple scoop cones canu make with 5 flavors of ice cream

Directrix (directrix):

By different, do you mean that the same flavor cannot be used for all three scoops of one cone?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

like different combinations

OpenStudy (anonymous):

like what make sense to me is 15..but when I write out the combination it doens make sense

Directrix (directrix):

What in the wording of the problem causes you to think that a cone that had 3 scoops of, say Vanilla, would not be allowed?

OpenStudy (anonymous):

right

Directrix (directrix):

Hey, let's look at a smaller problem. What if we were making two-scoop cones using the flavors of Vanilla and Chocolate. Then, would we have: VV, CC, VC? Would CV be considered different from VC?

OpenStudy (anonymous):

it says a local shop has 5 flavors of ice cream, how many different triple scoop cones can they make?

OpenStudy (anonymous):

no CV is the same as VC

Directrix (directrix):

What if we break this into 3 situations. Case I: The 3 scoops are all alike. There are 5 ways to do this. (Example: CCC, VVV, and so on for the 5 flavors). @vmami0204 Before we go to Case II, do you follow Case One?

Directrix (directrix):

Case 2: all three scoops are different The number of ways to choose three different flavors from a set of 5 flavors is "five choose three". C(5,3) = 10 (Use the combination formula.)

Directrix (directrix):

Case Three: two scoops the same, one scoop different That would be 5 choices for two alike flavors with 4 choices left for the third scoop. The number of possibilities is 5*4 = 20

Directrix (directrix):

In answer to the posted question: how many different triple scoop cones can u make with 5 flavors of ice cream ---------------- The sum of the case possibilities: 5 + 20 + 10 = 35 total. @vmami0204 Check my logic and my work and determine if you agree. Thanks.